consider a hypothetical economy where autonomous C ,autonomous I,autonomous G,autonomous TR ,mpc and t iss given
calculate the level of autonomuous expenditure,equilibrium income,consumption ,total taxes and multiplier
1
Expert's answer
2017-12-06T11:09:07-0500
If we know autonomous C (C0), autonomous I (I0), autonomous G (G0), autonomous TR (TR0), mpc and t, then:
Autonomous expenditure AE0 = C0 + I0 + G0,
equilibrium income Y = C + I + G,
consumption C = C0 + mpc*Y,
total taxes T = t*Y,
multiplier m = 1/mpc.